OpenStudy (anonymous):

As incentive for Gretchen to save the two thousand five hundred dollars she recently inherited, her parents tell her that they will pay her ten percent simple annual interest on the amount that she saves. How much money will Gretchen have after 5 years if she saves all of the money she inherited? two thousand six hundred forty dollars two thousand nine hundred fifty-five dollars three thousand fifty-six dollars three thousand seven hundred fifty dollars

OpenStudy (anonymous):

d

OpenStudy (uri):

Why don't you try doing it yourself rather than asking people for it. Did you try understanding the question? Even read it? @_lilb_to_smooth_4_U_

OpenStudy (anonymous):

explanation?

OpenStudy (uri):

@A.K.C.15 You're not allowed to give direct answers. http://openstudy.com/code-of-conduct

OpenStudy (anonymous):

@_lilb_to_smooth_4_U_ : Haven't I gave you so much formulas and reasoning before...Try to apply them

OpenStudy (anonymous):

\[I = \frac{ P \times R \times T }{ 100 }\]...Plug the values and calculate it and answer it

OpenStudy (anonymous):

well sorry @uri I'm new to this